# Geodesic Distance Between Graphs: A Spectral Metric for Assessing the Stability of Graph Neural Networks

### v1

@@ 4.2 Phase 2: Computing Geodesic Distances Between Graph Laplacians
The GGD metric can be formally defined as the infimum length of geodesics connecting two data points in the Riemannian manifold formed by the cone of the modified graph Laplacian matrices Lim et al. [2019]. This distance metric can be imagined as a matrix representation of the geometric distance $|\log(a/b)|$ between two positive numbers $a,b$ Bonnabel and Sepulchre [2010], Shamai and Kimmel [2017], Owen and Provan [2010].
@@
$GGD(G_{1},G_{2})=\left[\sum_{i=1}^{n}\log^{2}(\lambda_{i}(\mathcal{L}_{1}^{-1}\mathcal{L}_{2}))\right]^{1/2}$ (10)
@@
4.3 Connection between GGD and Graphs’ Structural Mismatches
Figure 2: The cut mismatch (for the node set $S$) between two simple graphs is $\frac{6}{2}=3$
For two graphs $G_{1}$ and $G_{2}$ that share the same node set $V$ (node correspondence is known in advance), let $L_{1}$ and $L_{2}$ be their respective Laplacian matrices. Consider a node subset denoted by $S$ and its complement denoted by $S^{\prime}$. We assign each node in $S$ a value of $1$, and each node in $S^{\prime}$ a value of 00. Then the node set $S$ can be defined as:
$S\stackrel{{\scriptstyle\text{def}}}{{=}}\{v\in V:x(v)=1\}$
The graph cut for node subset $S$, corresponding to the number of edges crossing $S$ and $S^{\prime}$ in graph $G_{1}$ can be computed by:
$\text{cut}_{G_{1}}(S,S^{\prime})=x^{T}L_{1}x$
As illustrated in Figure 2, the node subset $S$ has six edges crossing the boundary in $G_{1}$ but only two in $G_{2}$. This cut mismatch can be related to the generalized eigenvalue of the matrix pencil $(L_{1},L_{2})$ using the Generalized Courant-Fischer Minimax Theorem Golub and Van Loan [2013], Feng [2020].
@@ 4.3 Connection between GGD and Graphs’ Structural Mismatches
$\lambda_{k}=\min_{\begin{subarray}{c}\operatorname{dim}(U)=k\\ U\perp\operatorname{null}\left(L_{2}\right)\end{subarray}}\max_{x\in U}\frac{x^{\top}L_{1}x}{x^{\top}L_{2}x}$ (11)
This theorem allows us to find the upper bound of the maximum cut mismatch between two graphs by computing the most dominant generalized eigenvalue via the following optimization problem Feng [2020]:
$\lambda_{\max}=\max_{\begin{subarray}{c}|x|\neq 0\\ x^{\top}\mathbb{1}=0\end{subarray}}\frac{x^{\top}L_{1}x}{x^{\top}L_{2}x}\geq\max_{\begin{subarray}{c}|x|\neq 0\\ x(v)\in\{0,1\}\end{subarray}}\frac{x^{\top}L_{1}x}{x^{\top}L_{2}x}=\max\frac{cut_{G_{1}}(S,S^{\prime})}{cut_{G_{2}}(S,S^{\prime})}$ (12)
Based on (12), we can infer that the dominant generalized eigenvalue is closely related to the most significant graph cut mismatch between graphs $G_{1}$ and $G_{2}$. As a result, $\lambda_{\max}=\lambda_{1}$ connects to the upper bound of the cut mismatch between graphs $G_{1}$ and $G_{2}$, whereas $\lambda_{\min}=\lambda_{n}$ connects to the upper bound of the cut mismatch between graphs $G_{2}$ and $G_{1}$. Similarly, the second-largest (smallest) eigenvalues correspond to the next most significant cut mismatches. It is also obvious that the generalized eigenvalues close to $1$ will correspond to the minimum cut mismatches between two graphs.
@@ 5 Computing GGDs for Graphs with Different Sizes
To calculate geodesic distances between SPD matrices of different sizes, prior studies have proposed a submatrix adaptation method Lim et al. [2019]. In this approach, a principle submatrix with the same size as the smaller matrix is obtained from the larger matrix Ye and Lim [2016], and subsequently used to calculate the GGD. Furthermore, this method can be extended to project the smaller matrix into a larger one with the same size as the larger matrix Lim et al. [2019]. While these methods are efficient for handling SPD matrices, for our application taking the submatrix of the modified Laplacian can lose important nodes/edges, compromising critical graph structural properties.
@@
In this work, we will leverage spectral graph coarsening to address the issue. Spectral graph coarsening is a widely adopted process Loukas [2019], Aghdaei and Feng [2022] for reducing graph sizes while preserving key spectral (structural) properties, such as the Laplacian eigenvalues/eigenvectors. Recent spectral graph coarsening methods aim to decompose an input graph into many distinct node clusters, so that a reduced graph can be formed by treating each node cluster as a new node, with a goal of assuring that the reduced graph will approximately retain the original graph’s structure Loukas [2019], Han et al. [2024], Aghdaei and Feng [2022]. Therefore, when computing GGDs for graphs of different sizes, we can first adopt spectral graph coarsening to transform the bigger graph into a smaller one, so that our framework in Section 4 can be subsequently utilized. However, existing state-of-the-art graph coarsening methods do not allow us to precisely control the size of the reduced graphs.
@@ 5.1 Our Approach: Spectral Graph Coarsening by Effective Resistances
In this work, we introduce a spectral graph coarsening method using effective-resistance clustering Aghdaei and Feng [2022]. Our approach starts with estimating the effective resistances of all edges in the original graph. We can also incorporate the difference between node feature (if available) vectors as an additional parameter. In the graph coarsening phase, our method will rank edges according to their resistance distances and only the top few edges with the smallest resistances will be coarsened into new nodes. This approach enables precise control over the size of the reduced graphs while preserving crucial structural properties, such as the eigenvalues/eigenvectors of the adjacency matrices, which are essential for the subsequent spectral graph matching step (Phase 1 in Section 4.1).
Consider a connected, weighted, undirected graph $G=(V,E,w)$ with $|V|=n$. The effective resistance between nodes $(p,q)\in V$ plays a crucial role in various graph analysis tasks including spectral sparsification algorithms Spielman and Teng [2011]. The effective resistance distances can be accurately computed using the equation:
$R_{eff}(p,q)=\sum\limits_{i=2}^{n}\frac{w_{p,q}(u_{i}^{\top}b_{pq})^{2}}{\sigma_{i}},$ (13)
where ${b_{p}}\in\mathbb{R}^{n}$ denote the standard basis vector with all zero entries except for the $p$-th entry being $1$, and ${b_{pq}}=b_{p}-b_{q}$. $u_{i}\in\mathbb{R}^{n}$ for $i=1,\dots,n$ denote the unit-length, mutually-orthogonal eigenvectors corresponding to Laplacian eigenvalues $\sigma_{i}$ for $i=1,\dots,n$.
@@
To address the computational complexity associated with directly computing eigenvalues and eigenvectors required for estimating edge effective resistances, we leverage a scalable framework for approximating the eigenvectors of the graph Laplacian matrix using the Krylov subspace Saad [2011]. Let $A$ denote the adjacency matrix of graph $G$, consider its order-$m$ Krylov subspace $\mathbf{K}_{m}(A,x)$ that is a vector space spanned by the vectors computed through power iterations $x,Ax,A^{2}x,\ldots,A^{m-1}x$ Liesen and Strakos [2013]. By enforcing orthogonality among the above vectors in the Krylov subspace, a new set of mutually orthogonal vectors of unit lengths can be constructed for approximating the original Laplacian eigenvectors in 13, which are denoted as $\tilde{u}_{1}$, $\tilde{u}_{2}$, $\ldots$, $\tilde{u}_{m}$. To estimate the effective resistance between two nodes $p$ and $q$, we can exploit the approximated eigenvectors:
$R_{eff}(p,q)\approx\sum\limits_{i=1}^{m}\frac{w_{p,q}(\tilde{u}^{\top}_{i}b_{pq})^{2}}{\tilde{u}^{\top}_{i}L\tilde{u}_{i}},$ (14)
@@ 6 GGD as a Distance Metric
Assuming that the graph matching problem can always find the exact correspondence between nodes, we prove that the GGD metric between any two (nonempty) graphs is a pseudometric that satisfies the following conditions:
@@
  • (Positivity) The distance between two distinct graphs is positive: $GGD(G_{1},G_{2})>0$.
@@ 7 Experiments
7.1 Stability Analysis of GNNS
Figure 3: Correlation between graph distance metrics and GNN’s outputs
To analyze the stability of GNN models Bronstein et al. [2017], Garg et al. [2020], Duvenaud et al. [2015], we conducted multiple experiments with the GGD and TMD metrics. GNNs typically operate by a message-passing mechanism Gilmer et al. [2017], where at each layer, nodes send their feature representations to their neighbors. The feature representation of each node is initialized to its original features and is updated by repeatedly aggregating incoming messages from neighbors. In our experiment, we relate GGD to the Graph Isomorphism Networks (GIN) Xu et al. [2019], one of the most widely applied and powerful GNNs, utilizing the MUTAG dataset Morris et al. [2020] as our reference graph dataset. The objective is to analyze the relationship between the input distance $GGD(G_{1},G_{2})$ and the distance between the output GIN vectors, $\|h(G_{1})-h(G_{2})\|$ for randomly selected pairs of graphs. The result is illustrated in Figure 3.
We observe a strong correlation between GGD and the output distance, as indicated by a high Pearson correlation coefficient. This finding implies the effectiveness of the proposed GGD metric for analyzing the stability of GNN models Chuang and Jegelka [2022]. To compare GGD with existing metrics, we repeat this experiment using TMD without considering node attributes (features). As shown in Figure 3, GGD demonstrates a better correlation with GIN outputs than the TMD metric across different levels. These findings indicate that when dealing with graphs without node features, GGD should be adopted for the stability analysis of graph learning models.
7.2 Graph Classification
Table 2: Classification accuracies for various models on graph datasets
Accuracy in percentage
Dataset MUTAG PC-3H SW-620H BZR
@@ 7.2 Graph Classification
GCNKipf and Welling [2017] 77$\pm$3.8 70.56 69.4 72.56
GINXu et al. [2019] 82.6$\pm$4.6 75.34 73.4 77.09
WWLTogninalli et al. [2019] 72.4$\pm$2.6 65.46 68.34 73.59
WL SubtreeShervashidze et al. [2011] 76$\pm$6.3 68.43 70.56 N/A
FGHTitouan et al. [2019] 88.33$\pm$5.6 61.77 59.3 53.66
We evaluate whether the GGD metric aligns with graph labels in graph classification tasks using datasets from TUDatasets Morris et al. [2020]. We employ a Support Vector Classifier (SVC) $(C=1)$ with an indefinite kernel $e^{-\gamma*GGD(G_{1},G_{2})}$, which can be viewed as a noisy observation of the true positive semidefinite kernel Luss and d’Aspremont [2007]. The parameter $\gamma$ is selected through cross-validation from the set {0.01, 0.05, 0.1}. For comparative analysis with existing methods, we include graph kernels based on graph subtrees: the WL subtree kernel Shervashidze et al. [2011]; and two widely adopted GNNs: graph isomorphism network (GIN) Xu et al. [2019] and graph convolutional networks (GCN) Kipf and Welling [2017].
@@ 7.3 Runtime Complexity Analysis and Comparison
Table 3: Runtime comparison for different distance metrics on various datasets.
@@
When comparing various graph distance metrics, a primary consideration is their computational complexity. Conventional approaches usually require intricate computations that frequently have cubic time or higher complexities. For our problem, the spectral graph matching step requires the eigenvalue decomposition of adjacency matrices and solving the linear assignment problem (LAP). Eigenvalue decomposition of an $n\times n$ matrix has a complexity of $O(n^{3})$ Borodin and Munro [1975], André et al. [2015], while solving the LAP using the Hungarian algorithm also has a runtime complexity of $O(n^{3})$. Similarly, calculating the generalized eigenvalue of two SPD matrices entails a cubic complexity. Consequently, the overall complexity of GGD calculation is $O(n^{3})$. On the other hand, TMD is an OT-based distance metric with a complexity of $O(n^{3}\log(n))$ Chuang and Jegelka [2022], Flamary et al. [2021]. Therefore, GGD exhibits slightly better (lower) runtime complexity than the TMD metric.
To evaluate runtime performance, we conduct extensive experiments to compare the runtime of computing TMD at various levels with GGD on both small graphs (MUTAG, BZR) and large graphs (PC-3H, SW-620H) collected from the TUDataset Morris et al. [2020]. Table 3 presents the average runtime (in seconds) for computing $100$ distances between different graphs obtained by repeating the experiment five times. The results demonstrate that GGD consistently outperforms TMD in terms of runtime across all datasets and scenarios, particularly when dealing with larger graphs that contain more nodes. The TMD metric computation usually requires more levels to effectively capture the entire graph structure. In such cases, GGD exhibits runtime performance approximately $6-9$ times faster than of TMD. Hence, we conclude that GGD is significantly more computationally efficient than TMD, especially when working with large graphs. More details about our experimental setup can be found in Appendix A.4.

### v2

@@ 4.2 Phase 2: Computing Geodesic Distances Between Graph Laplacians
The GGD metric can be formally defined as the infimum length of geodesics connecting two data points in the Riemannian manifold formed by the cone of the modified graph Laplacian matrices (Lim et al., 2019). This distance metric can be imagined as a matrix representation of the geometric distance $|\log(a/b)|$ between two positive numbers $a,b$ (Bonnabel & Sepulchre, 2010; Shamai & Kimmel, 2017; Owen & Provan, 2010).
@@
$GGD(G_{1},G_{2})=\left[\sum_{i=1}^{n}\log^{2}(\lambda_{i}(\mathcal{L}_{1}^{-1}\mathcal{L}_{2}))\right]^{1/2},$ (10)
@@
The above GGD formulation for computing distances between SPD matrices is based on an Affine-Invariant Riemannian Metric (AIRM) (Lim et al., 2019), while another well-known metric, the Log-Euclidean Riemannian Metric (LERM) (Ilea et al., 2018; Thanwerdas & Pennec, 2023; Chen et al., 2024) is also discussed in Appendix A.7 of the supplementary section.
4.3 Connection between GGD and Graph Structural Mismatches
Consider two graphs, $G_{1}$ and $G_{2}$, that have the same node set $V$, with a known correspondence between their nodes. Let $L_{1}$ and $L_{2}$ be the Laplacian matrices of these graphs, respectively. Suppose we take a subset of nodes, denoted by $S$ and its complement, $S^{\prime}$. We assign the value $1$ to the nodes in $S$ and the value 00 to those in $S^{\prime}$. This defines the set $S$ as:
$S\stackrel{{\scriptstyle\text{def}}}{{=}}\{v\in V:x(v)=1\}.$
Figure 2: The cut mismatch (for the node set $S$) between two simple graphs is $\frac{6}{2}=3$.
For graph $G_{1}$, the cut for the node subset $S$ (which is the number of edges that cross between $S$ and $S^{\prime}$) can be computed as:
$\text{cut}_{G_{1}}(S,S^{\prime})=x^{T}L_{1}x.$
As shown in Figure 2, for node subset $S$ six edges cross between $S$ and $S^{\prime}$ in graph $G_{1}$, whereas only two edges cross in graph $G_{2}$. This difference in edge counts between the two graphs is referred to as a cut mismatch. The relationship between this cut mismatch and the generalized eigenvalue problem for the matrix pair $(L_{1},L_{2})$ can be formalized using the Generalized Courant-Fischer Minimax Theorem (Golub & Van Loan, 2013; Feng, 2020).
@@ 4.3 Connection between GGD and Graph Structural Mismatches
$\lambda_{k}=\min_{\begin{subarray}{c}\operatorname{dim}(U)=k\\ U\perp\operatorname{null}\left(L_{2}\right)\end{subarray}}\max_{x\in U}\frac{x^{\top}L_{1}x}{x^{\top}L_{2}x}.$ (11)
This theorem provides a method to bound the maximum cut mismatch between two graphs by calculating the largest generalized eigenvalue. Specifically, we can use the following optimization problem to compute the dominant eigenvalue $\lambda_{max}$ (Feng, 2020):
$\lambda_{\max}=\max_{\begin{subarray}{c}|x|\neq 0\\ x^{\top}\mathbb{1}=0\end{subarray}}\frac{x^{\top}L_{1}x}{x^{\top}L_{2}x}\geq\max_{\begin{subarray}{c}|x|\neq 0\\ x(v)\in\{0,1\}\end{subarray}}\frac{x^{\top}L_{1}x}{x^{\top}L_{2}x}=\max\frac{cut_{G_{1}}(S,S^{\prime})}{cut_{G_{2}}(S,S^{\prime})}.$ (12)
From equation (12), we can see that the dominant generalized eigenvalue $\lambda_{max}$ corresponds to the most significant cut mismatch between $G_{1}$ and $G_{2}$. In particular, $\lambda_{1}=\lambda_{max}$ sets an upper bound on the cut mismatch between $G_{1}$ and $G_{2}$, while $\lambda_{n}=\lambda_{min}$ defines the upper bound of the mismatch in the reverse direction, between $G_{2}$ and $G_{1}$.
@@ 5 Computing GGDs for Graphs with Different Sizes
To calculate geodesic distances between SPD matrices of different sizes, prior studies have proposed a submatrix adaptation method (Lim et al., 2019). In this approach, a principle submatrix with the same size as the smaller matrix is obtained from the larger matrix (Ye & Lim, 2016), and subsequently used to calculate the GGD. Furthermore, this method can be extended to project the smaller matrix into a larger one with the same size as the larger matrix (Lim et al., 2019). While these methods are efficient for handling SPD matrices, for our application taking the submatrix of the modified Laplacian can lose important nodes/edges, compromising critical graph structural properties.
@@
In this work, we will leverage spectral graph coarsening to address the issue. Spectral graph coarsening is a widely adopted process (Loukas, 2019; Aghdaei & Feng, 2022) for reducing graph sizes while preserving key spectral (structural) properties, such as the Laplacian eigenvalues/eigenvectors. Recent spectral graph coarsening methods aim to decompose an input graph into many distinct node clusters, so that a reduced graph can be formed by treating each node cluster as a new node, with a goal of assuring that the reduced graph will approximately retain the original graph’s structure (Loukas, 2019; Han et al., 2024; Aghdaei & Feng, 2022). Therefore, when computing GGDs for graphs of different sizes, we can first adopt spectral graph coarsening to transform the bigger graph into a smaller one, so that our framework in Section 4 can be subsequently utilized. However, existing state-of-the-art graph coarsening methods do not allow us to precisely control the size of the reduced graphs.
@@ 5.1 Our Approach: Spectral Graph Coarsening by Effective Resistances
In this work, we introduce a spectral graph coarsening method using effective-resistance clustering (Aghdaei & Feng, 2022). Our approach starts with estimating the effective resistances of all edges in the original graph. We can also incorporate the difference between node features (if available) as an additional parameter. In the graph coarsening phase, our method will rank edges according to their resistance distances and only the top few edges with the smallest resistances will be coarsened into new nodes. This approach enables precise control over the size of the reduced graphs while preserving crucial structural properties, such as the eigenvalues/eigenvectors of the adjacency matrices, which are essential for the subsequent spectral graph matching step (Phase 1 in Section 4.1).
Consider a connected, weighted, undirected graph $G=(V,E,w)$ with $|V|=n$. The effective resistance between nodes $(p,q)\in V$ plays a crucial role in various graph analysis tasks including spectral sparsification algorithms (Spielman & Teng, 2011). The effective resistance distances can be accurately computed using the equation:
$R_{eff}(p,q)=\sum\limits_{i=2}^{n}\frac{(u_{i}^{\top}b_{pq})^{2}}{\sigma_{i}},$ (13)
where ${b_{p}}\in\mathbb{R}^{n}$ denote the standard basis vector with all zero entries except for the $p$-th entry being $1$, and ${b_{pq}}=b_{p}-b_{q}$. $u_{i}\in\mathbb{R}^{n}$ for $i=1,\dots,n$ denote the unit-length, mutually-orthogonal eigenvectors corresponding to Laplacian eigenvalues $\sigma_{i}$ for $i=1,\dots,n$. A brief background on effective resistance is provided in Appendix A.3.
@@
To address the computational complexity associated with directly computing eigenvalues and eigenvectors required for estimating edge effective resistances, we leverage a scalable framework for approximating the eigenvectors of the graph Laplacian matrix using the Krylov subspace (Saad, 2011). Let $A$ denote the adjacency matrix of a graph $G$, consider its order-$m$ Krylov subspace $\mathbf{K}_{m}(A,x)$ that is a vector space spanned by the vectors computed through power iterations $x,Ax,A^{2}x,\ldots,A^{m-1}x$ (Liesen & Strakoš, 2012). By enforcing orthogonality among the above vectors in the Krylov subspace, a new set of mutually orthogonal vectors of unit lengths can be constructed for approximating the original Laplacian eigenvectors in 13, which are denoted as $\tilde{u}_{1}$, $\tilde{u}_{2}$, $\ldots$, $\tilde{u}_{m}$. To estimate the effective resistance between two nodes $p$ and $q$, we can exploit the approximated eigenvectors:
$R_{eff}(p,q)\approx\sum\limits_{i=1}^{m}\frac{(\tilde{u}^{\top}_{i}b_{pq})^{2}}{\tilde{u}^{\top}_{i}L\tilde{u}_{i}}+\alpha\|f_{p}-f_{q}\|,$ (14)
@@ 6 GGD as a Distance Metric
Assuming the graph matching problem can always find the exact correspondence between nodes, we prove that the GGD metric (based on AIRM) between any two nonempty graphs is a metric that satisfies the following conditions:
@@
  • (Positivity) The distance between two distinct graphs is positive: $GGD(G_{1},G_{2})\geq 0$.
@@ 7 Experiments
7.1 Application of GGDs in GNN Stability Analysis
Figure 3: Correlation between graph distance metrics and GNN model outputs.
To analyze the stability of GNN models (Bronstein et al., 2017; Garg et al., 2020; Duvenaud et al., 2015), we conducted multiple experiments with the GGD and TMD metrics. GNNs typically operate by a message-passing mechanism (Gilmer et al., 2017), where at each layer, nodes send their feature representations to their neighbors. The feature representation of each node is initialized to its original features and is updated by repeatedly aggregating incoming messages from neighbors. In our experiment, we relate GGD to the Graph Isomorphism Networks (GIN) (Xu et al., 2019), one of the most widely applied and powerful GNNs, utilizing the MUTAG dataset (Morris et al., 2020) as our reference graph dataset. The objective is to analyze the relationship between the input distance $GGD(G_{1},G_{2})$ and the distance between the output GIN vectors, $\|h(G_{1})-h(G_{2})\|$ for randomly selected pairs of graphs.
We employed a three-layer GIN network as described in (Xu et al., 2019). This network uses GIN convolutional layers to update tensors of nodes based on their neighboring nodes and then aggregates those outputs in a vector representation, followed by linear layers for classification tasks. Thus it outputs a single vector $h(G)$ for the entire graph $G$. The result is illustrated in Figure 3.
We observe a strong correlation between GGD and the output distance, as indicated by a high Pearson correlation coefficient. This finding implies the effectiveness of the proposed GGD metric for analyzing the stability of GNN models (Chuang & Jegelka, 2022). To compare GGD with existing metrics, we repeat this experiment using TMD without considering node attributes (features). As shown in Figure 3, GGD demonstrates a better correlation with GIN outputs than the TMD metric across different levels. These findings indicate that when dealing with graphs without node features, GGD should be adopted for the stability analysis of graph learning models.
7.2 Application of GGDs in Graph Classification Tasks
Table 2: Classification accuracies for various models on graph datasets.
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = c=4, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell4-122 = c, cell4-123 = c, cell4-124 = c, cell4-125 = c, hline1,3,13 = -, hline2 = 2-5, Dataset & Accuracy in percentage
MUTAG PC-3H SW-620H BZR
@@ 7.2 Application of GGDs in Graph Classification Tasks
GCN(Kipf & Welling, 2017) 77$\pm$3.8 70.56 69.4 72.56
GIN(Xu et al., 2019) 82.6$\pm$4.6 75.34 73.4 77.09
WWL(Togninalli et al., 2019) 72.4$\pm$2.6 65.46 68.34 73.59
WL Subtree(Shervashidze et al., 2011) 76$\pm$6.3 68.43 70.56 N/A
FGH(Titouan et al., 2019) 88.33$\pm$5.6 61.77 59.3 53.66
We evaluate whether the GGD metric aligns with graph labels in graph classification tasks using datasets from TUDatasets (Morris et al., 2020). We employ a Support Vector Classifier (SVC) $(C=1)$ with an indefinite kernel $e^{-\gamma*GGD(G_{1},G_{2})}$, which can be viewed as a noisy observation of the true positive semidefinite kernel (Luss & d’Aspremont, 2007). The parameter $\gamma$ is selected through cross-validation from the set {0.01, 0.05, 0.1}. For comparative analysis with existing methods, we include graph kernels based on graph subtrees: the WL subtree kernel (Shervashidze et al., 2011); and two widely adopted GNNs: graph isomorphism network (GIN) (Xu et al., 2019) and graph convolutional networks (GCN) (Kipf & Welling, 2017).
@@ 7.3 Runtime Complexity Analysis and Comparison
Table 3: Runtime comparison for different distance metrics on graph datasets.
@@
When comparing various graph distance metrics, a primary consideration is their computational complexity. Conventional approaches usually require intricate computations that frequently have cubic time or higher complexities. For our problem, the spectral graph matching step requires the eigenvalue decomposition of adjacency matrices and solving the linear assignment problem (LAP). Eigenvalue decomposition of an $n\times n$ matrix has a complexity of $O(n^{3})$ (Borodin & Munro, 1975; Flamary et al., 2021), while solving the LAP using the Hungarian algorithm also has a runtime complexity of $O(n^{3})$. Similarly, calculating the generalized eigenvalue of two SPD matrices entails a cubic complexity. Consequently, the overall complexity of GGD calculation is $O(n^{3})$. On the other hand, TMD is an OT-based distance metric with a complexity of $O(n^{3}\log(n))$ (Chuang & Jegelka, 2022; Flamary et al., 2021). Therefore, GGD exhibits slightly better (lower) runtime complexity than the TMD metric.
To evaluate runtime performance, we conduct extensive experiments to compare the runtime of computing TMD at various levels with GGD on both small graphs (MUTAG, BZR) and large graphs (PC-3H, SW-620H) collected from the TUDataset (Morris et al., 2020). Table 3 presents the average runtime (in seconds) for computing $100$ distances between different graphs obtained by repeating the experiment five times. The results demonstrate that GGD consistently outperforms TMD in terms of runtime across all datasets and scenarios, particularly when dealing with larger graphs that contain more nodes. The TMD metric computation usually requires more levels to effectively capture the entire graph structure. In such cases, GGD exhibits runtime performance approximately $6-9$ times faster than of TMD. Hence, we conclude that GGD is significantly more computationally efficient than TMD, especially when working with large graphs. More details about our experimental setup can be found in Appendix A.10.